A random sample of 10 chocolate energy bars of a certain brand has, on average, 230 calories per bar, with a standard deviation of 15 calories. Construct a 99% confidence interval for the true mean calorie content of this brand of energy bar. Assume that the distribution of the calorie content is approximately normal.

1
Expert's answer
2022-04-29T08:55:22-0400

The critical value for "\alpha = 0.01" and "df = n-1 = 9" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 3.249823."

The corresponding confidence interval is computed as shown below:


"CI=(\\bar{X}-t_c\\times\\dfrac{s}{\\sqrt{n}},\t\\bar{X}+t_c\\times\\dfrac{s}{\\sqrt{n}})"

"=(230-3.249823\times\dfrac{15}{\sqrt{10}},"

"230+3.249823\times\dfrac{15}{\sqrt{10}})"

"=(214.5847,245.4153)"

Therefore, based on the data provided, the 99% confidence interval for the population mean is "214.5847<\mu<245.4153," which indicates that we are 99% confident that the true population mean "\u03bc"  is contained by the interval "(214.5847,245.4153)."